OpenStudy (anonymous):

What is the slope of the graph of 2x – 6y = 9?

OpenStudy (anonymous):

Y=mx+b

OpenStudy (anonymous):

M is the slope

OpenStudy (anonymous):

oh sry a.1/3 b.-3 c.-1/3 d.3 B. C. D.

OpenStudy (lncognlto):

As @ayeshaafzal221 says, when the equation is in the form of y = mx + b, M is the slope. So put the equation into this form and what do you get?

OpenStudy (anonymous):

but im not shure what m x and b means

OpenStudy (anonymous):

????

OpenStudy (anonymous):

??????????????????

OpenStudy (lncognlto):

b is a constant where the graph cuts the y-axis when x = 0, and m is the slope of the graph.

OpenStudy (anonymous):

okay, but what would i do then?

OpenStudy (anonymous):

im sry i really have no idea how to do this :/

OpenStudy (skullpatrol):

$$\Huge 2x – 6y = 9$$ $$ \Huge 2x – 6y +6y= 9+6y$$ $$ \Huge 2x = 9+6y$$ $$ \Huge 2x-9 = 9+6y-9$$ $$ \Huge 2x-9 = 6y$$ $$ \Huge \frac{1}{6} ( 2x-9) = \frac{1}{6} 6y$$ $$ \Huge \frac{1}{3} x- \frac{3}{2} = y $$

OpenStudy (anonymous):

tahnks you so much!!!!!!!!!!

OpenStudy (skullpatrol):

Thanks for asking :-)